A small block of mass m is kept at the left end of a larger block of mass M and length L. The system can slide on a horizontal road.The system is started towards right with an initial velocity v.The friction coeffecient between the road and the bigger block is x and that between the two blocks is x/2.Find the time elapsed before the smaller block separate from the bigger block.

Grade:9

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

Let a1 and a2 be the accelerations of ma and M respectively.
Here, a1 > a2 so that m moves on M
it would be easiest if we see this in relative frame of M
see in relative frame
relative distance to be covered is l wrt to M
so we find a rel to M =a1-a2
and 1/2 a*t^2 = l
and for the force lower body M will have two force friction from the box below =-k(M+m)g
and fromt the upper body -k/2(mg)
a2=-(kmg-2kMg)/2M
a1=-k/2g
take the difference
-kg(M+m)/2M
now apply the kinematics formulae
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Ankit Jaiswal
8 years ago
Let a1 and a2 be the accelerations of ma and M respectively.
Here, a1 > a2 so that m moves on M
it would be easiest if we see this in relative frame of M
see in relative frame

relative distance to be covered is l wrt to M
so we find a rel to M =a1-a2
and 1/2 a*t^2 = l

and for the force lower body M will have two force friction from the box below =-k(M+m)g
and fromt the upper body -k/2(mg)
a2=-(kmg-2kMg)/2M
a1=-k/2g
take the difference
-kg(M+m)/2M

now apply the kinematics formulae